there is 10 liters of water and 10% of it is acid. how much more liters of water do you need to add for it to become 4% acid?

Answer by stanbon(75887)

there is 10 liters of water and 10% of it is acid. how much more liters of water do you need to add for it to become 4% acid?
----------------------------
Equation:
acid + acid = acid
0.10*10 + 0*x = 0.04(x+10)
Multiply thru by 100 to get:
10*10 = 4x+40
4x = 60
x = 15 liters (amt. of water you need to add to get the 4% mixture)
=================
Cheers,
Stan H.